# Learning Visual Hierarchies with Hyperbolic Embeddings

### v1

@@ 5.1 Part-based Image Retrieval
In Table 1, we show the top-$K{=}\{5,10,50,100\}$ retrieval accuracy for child-to-parent and parent-to-child. We notice a significant and consistent improvement with the fine-tuned models, in particular the hyperbolic model, across all metrics and model variants. This highlights the relevance of our fine-tuning set-up and the advantages of learning hierarchical part-based image embeddings in hyperbolic space.
Parent-to-child retrieval results show that the performance of fine-tuned models remains stable, while the zero-shot performance declines quickly. This trend likely occurs because zero-shot CLIP and MoCo-v2 rely primarily on visual similarity for retrieval; thus, their performance diminishes when the dataset lacks a sufficient number of visually similar objects for a given query. In contrast, after fine-tuning, the model learns the hierarchical structure of objects, enabling it to retrieve relevant objects even when they have substantially different appearances.
Vision Encoder Model Top-5 Top-10 Top-50 Top-100
@@
Hyp 73.37 72.59 69.84 68.62
Euc 67.83 68.37 67.12 66.04
CLIP ViT ZeroShot 26.73 25.95 23.69 22.70
Hyp 55.53 55.31 51.53 49.51
Euc 54.51 54.16 51.56 49.53
MoCo-v2 ZeroShot 17.23 16.86 16.33 16.07
@@
Hyp 66.02 66.63 66.50 65.91
Euc 65.38 65.70 66.01 65.79
CLIP ViT ZeroShot 47.52 46.60 43.50 42.08
Hyp 52.01 51.64 50.83 50.51
Euc 47.11 46.86 46.95 47.22
MoCo-v2 ZeroShot 18.49 18.37 17.73 17.45
Table 1: Part-based same-class image retrieval evaluation. For child-to-parent image retrieval, the retrieved parent must contain the object class of the query child. For parent-to-child image retrieval, the retrieved child must match a class within the parent.
[…]
@@
Hyp 47.28 74.06 16.36
Euc 46.32 72.96 17.00
K=150,000 ZeroShot 36.45 51.42 24.30
Hyp 42.37 82.34 19.27
Euc 41.47 80.98 20.00
K=200,000 ZeroShot 35.61 64.27 29.94
Hyp 38.37 88.79 22.21
Euc 38.09 87.97 22.45
K=250,000 ZeroShot 35.21 76.91 25.29
@@
Hyp 46.70 56.09 12.41
Euc 45.70 55.81 13.72
K=150,000 ZeroShot 38.34 48.07 23.41
Hyp 43.81 71.61 15.29
Euc 43.07 71.86 16.83
K=200,000 ZeroShot 37.18 62.56 24.63
Hyp 40.41 83.13 19.09
Euc 39.99 83.42 20.17
K=250,000 ZeroShot 36.24 76.77 25.65
Table 2: Hierarchical evaluation of parent-to-child image retrieval. OT distance is defined in Sec. 3.3.
Hierarchical Retrieval.
Table 2 evaluates the hierarchical structure of the embedding space by retrieving a large number of child images from parent images. We check the alignment between the retrieved distribution and the underlying hierarchical distribution of the test set. Good alignment is a desirable property for fine-grained retrieval as the retrieved set should capture the hierarchies present in the data distribution. The alignment is measured by the optimal transport (Wasserstein distance) between two distributions, with a smaller distance indicating a closer match (see Sec. 3.3). Precision measures the percentage of the retrieved images that are correctly positioned on the lower level of the ground truth hierarchy tree, while recall measures the percentage of ground truth images that are successfully retrieved.
As shown in Table 2, the fine-tuned models in Euclidean and hyperbolic space are significantly improved over the zero-shot CLIP and MoCo-v2 models. The model trained in hyperbolic space better represents the hierarchical distribution of the test set compared to the Euclidean model, outperforming the latter in all cases for precision, and in 4 out of 6 instances for the recall.
@@
To evaluate the emergent behaviors of models trained in hyperbolic versus Euclidean spaces, we trained the CLIP ViT model solely on hierarchical part-based entailment data within individual images (entailment pairs with high visual similarity), omitting any cross-image image-to-bounding-box samples.
Different from Top-K retrieval, we compute hierarchical precision and recall metrics as shown in Fig. 5, by applying progressively increasing angle thresholds for $\beta_{1}$ (range 00 to $\pi$) in the fine-tuned models and cosine similarity thresholds (range 0 to 1) for the zero-shot CLIP model to generate the Precision-Recall (PR) curve. As shown, the model fine-tuned in hyperbolic space without cross-image samples outperforms the Euclidean model by a large margin and is even better than the model trained with additional cross-image samples. This clearly indicates the hyperbolic model learns the structure implicitly. When both models are trained with cross-image samples, the hyperbolic model still exhibits better performance than the Euclidean model, however the performance gap is reduced.
Figure 4: Example of parent-to-child retrieval using CLIP ViT zero-shot and fine-tuned model. Results are ordered by embedding norms. Our model retrieves images matching the predefined scene-object-part hierarchy, beyond visual similarity. Best viewed when zoomed in.
@@
In Fig. 4, we show examples of parent-to-child retrieval, with results ordered by increasing embedding norm. The hyperbolic model effectively learns the predefined scene-object-part hierarchy, placing high-level objects near the origin. It captures diverse, visually different objects that might appear in the scene, even if absent from the query image. In contrast, the zero-shot model struggles to link large scenes with smaller objects.
Figure 5: Precision-Recall curves of zero-shot and fine-tuned CLIP ViT models. Dotted lines show models trained solely on hierarchical part-based entailment data within the same images; solid lines represent models trained with additional cross-image scene-to-object samples. Angle or cosine similarity threshold values are marked by text.
5.2 Zero-shot Image Classification
We evaluate the generalization properties of the fine-tuned CLIP model on zero-shot image classification task among 10 popular datasets as shown in Table 3. Consistent with the retrieval results, our model shows significant improvement in zero-shot image classification accuracy on 8 of 10 datasets, with the hyperbolic model consistently outperforming the Euclidean model. This indicates that the learned hyperbolic embeddings effectively capture the underlying structure and relationships within the data, allowing for stable and reliable predictions. This also suggests that the latent space is both well-conditioned and generalizable, as we observe consistent performance across diverse datasets.
@@ 5.2 Zero-shot Image Classification
Hyp 84.1 49.7 38.3 79.7 94.7 33.3 42.3 20.5 34.4 36.9
Euc 80.2 44.8 31.8 78.3 93.1 29.2 23.8 4.8 11.6 7.0
ZeroShot 55.1 28.7 39.7 62.2 63.5 11.8 15.0 14.3 31.6 40.0
Table 3: Zero-shot image classification evaluation of CLIP ViT (Top-5 majority vote). Accuracy is in percentage.
@@ 6 Conclusion
In this work, we introduce a new learning paradigm that effectively encodes user-defined multi-level visual hierarchies in hyperbolic space without requiring explicit hierarchical labels. We present a concrete example of defining a multi-level part-based image hierarchy using object-level annotations and propose a contrastive loss in hyperbolic space to enforce pairwise entailment relationships. Additionally, we introduce new evaluation metrics for hierarchical image retrieval. Our experiments demonstrate our model effectively learns the predefined image hierarchy and goes beyond visual similarity.
@@ Appendix A Hierarchy Tree
We introduce a general method to generate dataset-specific ground truth hierarchy trees based on data statistics. Following the approach outlined in Sec. 3.1 of the main paper, we begin by identifying bounding box pairs with substantial overlap. These pairs are then filtered based on two criteria: a frequency threshold and a proportion threshold. For each pair, we record the frequency of occurrences (e.g., bicycle-to-wheel relationships) and calculate the proportion of instances where a child class appears within a parent class (e.g., the percentage of bicycle bounding boxes containing a wheel bounding box). Only pairs meeting both criteria, frequent occurrence and consistent labeling, are preserved. We choose $\textit{frequency}=50$ and $\textit{proportion}=10\%$ in our experiments.
@@ B.1 Hyperparameters and training details
We employ the AdamW optimizer with parameters $(\beta_{1},\beta_{2})=(0.9,0.999)$ and a learning rate of $2\times 10^{-5}$. Training was conducted using 8 $\times$ A10G Nvidia GPUs. For each model, we used the largest batch size that fit in memory: CLIP ViT was trained with a total effective batch size of 320, and MoCo-v2 with a total effective batch size of 800. Each model was fine-tuned for a single epoch on HierOpenImages dataset, taking approximately 26 hours for CLIP ViT and 18 hours for MoCo-v2.
@@ B.2 Part-based Image Retrieval
The HierOpenImages test set contains approximately 1 million full images and 3 million bounding box images. During part-based image retrieval evaluation, to limit the number of object classes within a bounding box image, we only evaluate on bounding box images that occupy 5%-30% of the full image area. This reduced the candidate full images and bounding box images to 0.33M and 0.6M, respectively.
For part-based image retrieval, we randomly select a subset of 10,000 full images and 10,000 bounding box images as query images, using the entire test set as candidates. The bounding box images in the query set span 339 object classes, while the full images are labeled with 516 classes. The top-50 class frequency distributions for both the query and candidate sets are shown in Fig. 7. Although the query and candidate set distributions are similar, the class distribution is highly imbalanced, highlighting the importance of hierarchical retrieval evaluation using combined precision-recall curves and OT distances (Sec. 3.3; see Fig. 5 and Table 2 in the main paper).
@@
Metric Model Top-5 Top-10 Top-50 Top-100
Child-to-Parent
Cos Sim. HCL 15.29 15.34 15.12 14.95
Hyp Dist. HCL 14.64 14.72 14.45 14.27
Hyp Angle Ours 55.53 55.31 51.53 49.51
Parent-to-Child
Cos Sim. HCL 17.10 16.90 16.45 16.23
Hyp Dist. HCL 16.72 16.52 16.20 16.00
Hyp Angle Ours 52.01 51.64 50.83 50.51
Table 4: Part-based same-class image retrieval evaluation on HCL [13] model and our model.
[…]
Cos Sim.
K=150,000 HCL 36.05 46.29 24.46
K=200,000 HCL 35.82 61.38 24.87
K=250,000 HCL 35.62 76.43 25.24
Hyp Dist.
K=150,000 HCL 35.60 45.39 25.10
K=200,000 HCL 35.32 60.22 25.63
K=250,000 HCL 35.15 75.23 25.96
Hyp Angle
K=150,000 Ours 46.70 56.09 12.41
K=200,000 Ours 43.81 71.61 15.29
K=250,000 Ours 40.41 83.13 19.09
Table 5: Hierarchical evaluation of parent-to-child image retrieval on HCL [13] model and our model. The OT distance is defined in Sec. 3.3 in the main paper.
@@
Table 6: Query and candidate set sizes for each dataset used in zero-shot image classification evaluation.
@@
B.3 Zero-shot Image Classification
@@
B.4 Comparison to recent hyperbolic image model.
We compared our method with the recent hyperbolic image model, HCL [13], on the part-based same-class and hierarchical image retrieval tasks (see Table 4-5). HCL is trained on the OpenImages dataset using scene-to-object relationships and demonstrates hierarchical structure in its embeddings, making it, to our knowledge, the most similar method to ours.
HCL learns scene-to-object relationships by minimizing the hyperbolic distance between scene images and their corresponding object bounding box images, while also maximizing the cosine similarity between slightly shifted and jittered crops of the same object. We evaluated its retrieval performance using both cosine similarity and hyperbolic distance metrics. We compare our MoCo-v2 hyperbolic model with HCL, as both use the same visual backbone. Instead of fine-tuning, HCL retrained MoCo-v2 on the OpenImages dataset. HCL underperformed compared to the pretrained MoCo-v2 model trained purely in Euclidean space (see Table 1-2 in the main paper), and performed significantly worse than our model on these tasks. These results indicate that while HCL is conceptually similar, it is not a suitable comparison algorithm for our task.
@@ Appendix C Ablation Studies
In this ablation study, we evaluate the impact of cross-image scene-to-object samples. We fine-tuned the CLIP ViT model solely on hierarchical part-based entailment data within individual images (entailment pairs with high visual similarity), excluding any cross-image image-to-bounding-box samples. Fig. 5 in the main paper shows that the hyperbolic model fine-tuned without cross-image samples significantly outperforms the Euclidean model and even surpasses the model trained with additional cross-image samples. This indicates that training in hyperbolic space enhances the model’s ability to recognize visually dissimilar entailment pairs.
In this supplementary material, we further evaluate these models on part-based same-class and hierarchical image retrieval tasks, as shown in Table. 7-8. The best results are highlighted in bold and the second-best results are shown in blue. The results clearly indicate that cross-image sampling improves image retrieval performance. Notably, the results align with Fig. 5 in the main paper, with the hyperbolic model trained without cross-image samples outperforming the Euclidean model trained with cross-image samples in 11 out of 17 cases.
@@
Hyp 73.37 72.59 69.84 68.62
Euc 67.83 68.37 67.12 66.04
Hyp 70.47 70.29 67.89 66.38
Euc 64.98 64.49 63.51 63.10
ZeroShot - 26.73 25.95 23.69 22.70
@@
Hyp 66.02 66.63 66.50 65.91
Euc 65.38 65.70 66.01 65.79
Hyp 63.00 63.78 63.65 63.44
Euc 60.33 60.73 61.43 61.38
ZeroShot - 47.52 46.60 43.50 42.08
Table 7: Part-based same-class image retrieval evaluation on fined-tuned and zero-shot CLIP ViT models. Cross-image ✓indicates models fine-tuned on entailment relationships both within and across images at the category level, while ✗ represents models fine-tuned without cross-image sampling. The evaluation setup is the same as Table 1 in the main paper. The best results are marked in bold, and the second-best results are in blue.
@@
Hyp 47.28 74.06 16.36
Euc 46.32 72.96 17.00
Hyp 46.87 73.52 16.95
Euc 44.93 71.34 18.43
K=150,000 ZeroShot - 36.45 51.42 24.30
Hyp 42.37 82.34 19.27
Euc 41.47 80.98 20.00
Hyp 41.94 81.58 19.77
Euc 40.02 78.89 21.25
K=200,000 ZeroShot - 35.61 64.27 29.94
Hyp 38.37 88.79 22.21
Euc 38.09 87.97 22.45
Hyp 38.00 88.10 22.55
Euc 36.87 86.04 23.53
K=250,000 ZeroShot - 35.21 76.91 25.29
Table 8: Hierarchical evaluation of parent-to-child image retrieval on fined-tuned and zero-shot CLIP ViT models. The OT distance is defined in Sec. 3.3 (main paper), and the evaluation setup follows Table 2 (main paper). The best results are marked in bold, and the second-best results are in blue.

### v2

@@ 5.1 Main Results
[…]
Vision Encoder Model Metrics Top-5 Top-10 Top-50 Top-100
@@
CLIP ViT CLIP Cos Sim. 26.73 25.95 23.69 22.70
CLIP-euc Euc Ang. 67.83 68.37 67.12 66.04
CLIP-hyp Hyp Ang. 73.37 72.59 69.84 68.62
HCL Cos Sim. 15.11 14.49 14.69 14.60
HCL Hyp Dist. 14.46 14.34 14.04 13.92
HCL Hyp Dist. 14.54 14.43 14.16 14.01
MoCo-v2 MoCo Cos Sim. 17.23 16.86 16.33 16.07
MoCo-euc Euc Ang. 54.51 54.16 51.56 49.53
MoCo-hyp Hyp Ang. 55.53 55.31 51.53 49.51
@@
CLIP ViT CLIP Cos Sim. 47.52 46.60 43.50 42.08
CLIP-euc Euc Ang. 65.38 65.70 66.01 65.79
CLIP-hyp Hyp Ang. 66.02 66.63 66.50 65.91
HCL Cos Sim. 16.08 15.93 15.84 15.74
HCL Hyp Dist. 16.16 16.04 15.60 15.45
HCL Hyp Dist. 17.07 16.57 15.97 15.69
MoCo-v2 MoCo Cos Sim. 18.49 18.37 17.73 17.45
MoCo-euc Euc Ang. 47.11 46.86 46.95 47.22
MoCo-hyp Hyp Ang. 52.01 51.64 50.83 50.51
Table 1: Part-based same-class image retrieval evaluation. For child-to-parent image retrieval, the retrieved parent must contain the object class of the query child. For parent-to-child image retrieval, the retrieved child must match a class within the parent. $\dagger$ denotes models fine-tuned on the HierOpenImages dataset. Our proposed method is shaded in purple.
Metrics Model Dist. Func. Top-150k Top-200k Top-250k
@@
Recall $\uparrow$ CLIP Cos Sim. 51.33 64.28 77.02
CLIP-euc Euc Ang. 72.96 80.98 87.97
CLIP-hyp Hyp Ang. 74.06 82.34 88.79
OT Distance $\downarrow$ CLIP Cos Sim. 23.81 24.60 25.03
CLIP-euc Euc Ang. 17.00 20.00 22.45
CLIP-hyp Hyp Ang. 16.36 19.27 22.21
@@
Recall $\uparrow$ HCL Cos Sim. 46.05 61.24 76.44
HCL Hyp Dist. 46.06 61.14 76.16
HCL Hyp Dist. 45.81 60.77 75.77
MoCo Cos Sim. 48.38 63.63 77.89
MoCo-euc Euc Ang. 55.81 71.86 83.42
MoCo-hyp Hyp Ang. 56.09 71.61 83.13
OT Distance $\downarrow$ HCL Cos Sim. 24.23 24.67 25.09
HCL Hyp Dist. 25.72 25.98 26.04
HCL Hyp Dist. 26.12 26.32 26.28
MoCo Cos Sim. 24.05 24.44 24.93
MoCo-euc Euc Ang. 13.72 16.83 20.17
MoCo-hyp Hyp Ang. 12.41 15.29 19.09
Table 2: Part-based hierarchical evaluation of parent-to-child image retrieval on HierOpenImages. Results are evaluated using the ground truth hierarchy tree and the hierarchical distribution of the test set. Smaller OT distance indicates better distribution alignment (ref. Sec. 3.3).
Hierarchical Retrieval via the Learned Latent Space Distribution.
Table 2 evaluates the hierarchical structure of the latent space by retrieving a large number of child images from parent images. We use recall to measure the percentage of ground truth images that are successfully retrieved. Moreover, we check the alignment between the retrieved distribution and the underlying hierarchical distribution of the full test set. Good distribution alignment is a desirable property for fine-grained retrieval as the retrieved set should capture the hierarchies present in the data distribution. We propose to measure distribution alignment using the optimal transport (Wasserstein distance), with a smaller distance indicating a closer match (see Sec. 3.3).
As shown in Table 2, our hyperbolic model better captures the hierarchical distribution of the test set compared to the Euclidean model, achieving better OT distance in all cases, and in 4 out of 6 instances for the recall.
All fine-tuned models using our angle-based entailment loss show significant improvement over the baseline models. Notably, even after fine-tuning, HCL [14] shows a decline in hierarchical retrieval performance, indicating its difficulty in learning the complex visual hierarchies in the training data.
@@
To compare the emergent behaviors of the hyperbolic and Euclidean models, we trained the CLIP ViT model solely on hierarchical part-based entailment data within individual images (entailment pairs with high visual similarity), omitting any cross-image image-to-bounding-box samples.
Fig. 5 shows the Precision-Recall (PR) curve with increasing $\beta_{1}$ angle thresholds (00 to $\pi$) for CLIP-hyp and CLIP-euc, and with cosine similarity thresholds (0 to 1) for the baseline CLIP. The proposed hyperbolic model trained without cross-image samples substantially outperforms the Euclidean model and even exceeds the Euclidean model with cross-image samples, indicating it implicitly learns the underlying structure. When cross-image samples are introduced, the hyperbolic model still leads, though the performance gap narrows.
Figure 4: Example of parent-to-child retrieval using CLIP ViT and our CLIP-hyp model. Results are ordered by ascending norms. Our model retrieves images matching the predefined scene-object-part hierarchy, placing high-level objects near the origin (e.g, harbor $\rightarrow$ boart parts), and grouping semantically related but visually distinct objects (e.g, microwave oven & kitchen hood).
@@
Following [9, 15, 40], we use parent-to-child image traversals with results ordered by increasing embedding norm, to illustrate the latent space, which: (1) aligns with the predefined scene-object-part hierarchy, placing high-level objects near the origin. (2). groups diverse objects under the same lower hierarchical branch, even if they are visually distinct (e.g., keyboard and monitor under the studio image). In contrast, CLIP struggles to link large scenes with smaller objects.
Figure 5: Precision-Recall curves of CLIP ViT models on hierarchical retrieval. Dotted lines show models trained only on hierarchical entailment data within the same images; solid lines represent models trained with additional cross-image scene-to-object samples. Angle or cosine similarity threshold values are marked by text.
5.2 Generalization Evaluation
We evaluate the generalization of our method through image retrieval on unseen datasets and hierarchies. Similar to Table 1-2, Table 3 presents part-based same-class and hierarchical retrieval performance on the out-of-domain LVIS dataset [17], which contains 1,203 fine-grained object classes, including many rare objects absent from the training HierOpenImages dataset. Table 4 evaluates image retrieval performance on 10 popular single-class datasets using top-5 majority voting. Results demonstrate that our visual hierarchical learning significantly enhances model generalization on unseen datasets and image hierarchies.
Model Dist. Func. Top-5 Top-10 Top-50 Top-100
Child-to-Parent
CLIP Cos Sim. 2.37 2.32 2.04 1.84
CLIP-euc Euc Ang. 19.36 18.64 16.14 14.52
CLIP-hyp Hyp Ang. 22.33 21.40 18.50 16.42
Parent-to-Child
CLIP Cos Sim. 8.57 7.77 5.70 4.64
CLIP-euc Euc Ang. 20.24 20.03 19.30 18.68
CLIP-hyp Hyp Ang. 22.41 22.29 21.70 21.10
Metrics Model Dist. Func. Top-10k Top-20k Top-30k
Recall $\uparrow$ CLIP Cos Sim. 18.30 32.42 45.42
CLIP-euc Euc Ang. 46.89 64.40 75.71
CLIP-hyp Hyp Ang. 47.57 64.71 75.72
OT Distance $\downarrow$ CLIP Cos Sim. 7.61 8.88 9.43
CLIP-euc Euc Ang. 5.74 7.39 8.43
CLIP-hyp Hyp Ang. 5.74 7.38 8.40
Table 3: Out-of-domain part-based image retrieval evaluation on the LVIS dataset: same-class (top) and part-based hierarchical retrieval (bottom).
@@ 5.2 Generalization Evaluation
CLIP 55.1 28.7 39.7 62.2 63.5 11.8 15.0 14.3 31.6 40.0
CLIP-euc 80.2 44.8 31.8 78.3 93.1 29.2 23.8 4.8 11.6 7.0
CLIP-hyp 84.1 49.7 38.3 79.7 94.7 33.3 42.3 20.5 34.4 36.9
Table 4: Unseen single-class image retrieval (image-only model). For each query image, retrieve the top-5 images based on cosine similarity, and predict the class by majority vote among these top images.
@@ 6 Conclusion
In this work, we introduce a new learning paradigm that effectively encodes user-defined visual hierarchies in hyperbolic space without requiring explicit hierarchical labels. We present a concrete example of defining a part-based multi-level complex image hierarchy using object-level annotations and propose a contrastive loss in hyperbolic space to enforce pairwise entailment relationships. Additionally, we introduce new evaluation metrics for hierarchical image retrieval. Our experiments demonstrate our model effectively learns the predefined image hierarchy and goes beyond visual similarity.
@@ Appendix A Hierarchy Tree
We introduce a general method to generate dataset-specific ground truth hierarchy trees based on data statistics. Following the approach outlined in Sec. 3.1 of the main paper, we begin by identifying bounding box pairs with substantial overlap. In this work, we define a bounding box pair when at least 80% of the smaller bounding box $b$ is contained within either the full image or a larger bounding box. We initially set the overlap threshold at 100% and empirically adjusted it by evaluating the hierarchy’s validity using text labels. For the OpenImages dataset, 80% was found to be the most suitable threshold. This is a design choice and can be adapted for other datasets.
These pairs are then filtered based on two criteria: a frequency threshold and a proportion threshold. For each pair, we record the frequency of occurrences (e.g., bicycle-to-wheel relationships) and calculate the proportion of instances where a child class appears within a parent class (e.g., the percentage of bicycle bounding boxes containing a wheel bounding box). Only pairs meeting both criteria, frequent occurrence and consistent labeling, are preserved. We choose $\textit{frequency}=50$ and $\textit{proportion}=10\%$ in our experiments.
@@ B.1 Hyperparameters and training details
We employ the AdamW optimizer with parameters $(\beta_{1},\beta_{2})=(0.9,0.999)$ and a learning rate of $2\times 10^{-5}$. Training was conducted using 8 $\times$ A10G Nvidia GPUs. For each model, we used the largest batch size that fit in memory: CLIP ViT was trained with a total effective batch size of 320, and MoCo-v2 with a total effective batch size of 800. Each model was fine-tuned for a single epoch on HierOpenImages dataset, taking approximately 26 hours for CLIP ViT and 18 hours for MoCo-v2. The embeddings are projected to dimension 128 in the final layer. The hyperbolic model has a learnable curvature parameter.
@@ B.2 Part-based Image Retrieval
The HierOpenImages test set contains approximately 1 million full images and 3 million bounding box images. During part-based image retrieval evaluation, we filter out bounding boxes for very small or large objects that occupy less than 5% or more than 30% of the full image area. This reduces the candidate full images to 59,691 and bounding box images to 330,063. For part-based image retrieval, we randomly select a subset of 10,000 full images and 10,000 bounding box images as query images, using the entire test set as candidates. The bounding box images in the query set span 339 object classes, while the full images are labeled with 516 classes. The top-50 class frequency distributions for both the query and candidate sets are shown in Fig. 7. Although the query and candidate set distributions are similar, the class distribution is highly imbalanced, highlighting the importance of hierarchical retrieval evaluation using combined precision-recall curves and OT distances (Sec. 3.3; see Fig. 5 and Table 2 in the main paper).
@@
Table 5: Query and candidate set sizes for each single-class unseen dataset.
@@
Figure 8: Example of our image retrieval interface built with Gradio [1]. The interface supports image selection/upload and retrieves results ranked by user-defined modes. Its modular design allows for easy integration of additional functionalities.
Image Retrieval Interface via Gradio.
We built our image retrieval interface using Gradio [1], as shown in Fig. 8. Input images can be selected from a linked image folder, where thumbnail images are displayed in an image gallery, or they can be directly uploaded by users. The retrieval results can be sorted by the hyperbolic angle relative to the input image or filtered using a user-defined threshold value, after which the results are ordered by their embedding norms. Additional functionalities can be easily integrated into the current pipeline.
B.3 Generalization Evaluation
LVIS Dataset.
We evaluate the generalization capability of our model on the out-of-domain LVIS dataset [17], which is designed for long-tail instance classification and segmentation. It has a highly imbalanced distribution of 1,203 object categories and contains 897 object categories that are absent from OpenImages [28]. The full list of these categories can be found in the appendix. Only display the first class of synonyms.
We construct the hierarchical evaluation set from the validation set of the LVIS dataset [17], which contains 19,809 images. Following the same process as constructing HierOpenImages, we evaluate bounding box images and corresponding full images that occupy 5%-30% of the full image area, reducing the candidate set to 14,716 full images and 76,255 bounding box images. We use the full reduced set for part-based image retrieval in Table 3 in the main paper.
[…]
Single-class Datasets.
@@ Appendix C Ablation Studies
In this ablation study, we evaluate the impact of cross-image scene-to-object samples. We fine-tuned the CLIP ViT model solely on hierarchical part-based entailment data within individual images (entailment pairs with high visual similarity), excluding any cross-image image-to-bounding-box samples. Fig. 5 in the main paper shows that the hyperbolic model CLIP-hyp fine-tuned without cross-image samples significantly outperforms the Euclidean model CLIP-euc and even surpasses the model trained with additional cross-image samples. This indicates that training in hyperbolic space enhances the model’s ability to recognize visually dissimilar entailment pairs.
In this supplementary material, we further evaluate these models on part-based same-class and hierarchical image retrieval tasks, as shown in Table. 6-7. The best results are highlighted in bold and the second-best results are shown in blue. The results clearly indicate that cross-image sampling improves image retrieval performance. Notably, the hierarchical retrieval results align with Fig. 5 in the main paper. Specifically, the hyperbolic model trained without cross-image samples outperforms the Euclidean model trained with cross-image samples in 5 out of 6 cases, as shown in Table 7.
@@
CLIP-hyp 73.37 72.59 69.84 68.62
CLIP-euc 67.83 68.37 67.12 66.04
CLIP-hyp 70.47 70.29 67.89 66.38
CLIP-euc 64.98 64.49 63.51 63.10
CLIP - 26.73 25.95 23.69 22.70
@@
CLIP-hyp 66.02 66.63 66.50 65.91
CLIP-euc 65.38 65.70 66.01 65.79
CLIP-hyp 63.00 63.78 63.65 63.44
CLIP-euc 60.33 60.73 61.43 61.38
CLIP - 47.52 46.60 43.50 42.08
Table 6: Part-based same-class image retrieval evaluation. Cross-image ✓indicates models fine-tuned on entailment relationships both within and across images at the category level, while ✗ represents models fine-tuned without cross-image sampling. The evaluation setup is the same as Table 1 in the main paper. The best results are marked in bold, and the second-best results are in blue.
Metrics Model Cross-Image Top-150k Top-200k Top-250k
CLIP-hyp 74.06 82.34 88.79
CLIP-euc 72.96 80.98 87.97
CLIP-hyp 73.52 81.58 88.10
CLIP-euc 71.34 78.89 86.04
@@
CLIP-hyp 16.36 19.27 22.21
CLIP-euc 17.00 20.00 22.45
CLIP-hyp 16.95 19.77 22.55
CLIP-euc 18.43 21.25 23.53
OT Distance $\downarrow$ CLIP - 23.81 24.60 25.03
Table 7: Part-based hierarchical evaluation of parent-to-child image retrieval. The OT distance is defined in Sec. 3.3 (main paper), and the evaluation setup follows Table 2 (main paper). The best results are marked in bold, and the second-best results are in blue.